OpenStudy (anonymous):

Solve the compound inequality. 2x>-10 and 9x<18. I don't know how to do this. Can someone help me with the steps?

OpenStudy (anonymous):

First I would isolate x. If you do that, then you get x>-5 and x<2, now you simply get an interval where x can bee any number. -5 < x < 2

OpenStudy (adilalvi):

2x>-10= 9x<18 2x-9x=18+10 -7x=28 -x=28/7 -x=4 now multiply both sides by - x=-4 answer